OpenStudy (anonymous):

A restaurant charges $8 for a large cheese pizza plus $1.50 per topping. Another restaurant charges $11 for a large cheese pizza plus $0.75 per topping. To find out how many toppings it would take for the pizzas to cost the same at both restaurants,

OpenStudy (anonymous):

Let N be the number of toppings. At first pizza place, the cost of pizza = 8 + 1.5N At second pizza place, the cost of pizza = 11 + 0.75N Solve for N when both are same.

OpenStudy (anonymous):

Do u get 8 or 10

OpenStudy (anonymous):

I got 4. 8 + 1.5N = 11 + 0.75N 1.5N - 0.75N = 11 - 8 0.75N = 3 N = 3/0.75 N = 4

OpenStudy (anonymous):

??

OpenStudy (anonymous):

No response. No medal. Wow.